Proxima Centauri is the closest star to the Earth. The angle theta is 0.762 seconds of arc, and the radius of the orbit of the Earth around the sun is 1.5 x 10^11 m. Calculate the distance to Proxima Centauri, in m.


1
Expert's answer
2021-04-13T06:39:25-0400

The distance is


"D=\frac{r}{\tan\theta}=\frac{1.5·10^{11}}{\tan0.762"}=4.06\u00b710^{16}\\text{ m}."